A 63.0kg sprinter starts a race with an acceleration of 4.20m/s square. What is the net external force on him?
If the sprinter from the previous problem accelerates at that rate for 20m, and then maintains that velocity for the remainder for the 100-m dash, what will be his time for the race?
1
Expert's answer
2019-07-01T09:27:03-0400
"F=ma=(63)(4.2)=265\ N"

We have:


"v^2=2as_1=2(4.2)(20)"

"v=12.96\frac{m}{s}"

"T=t_1+t_2=\frac{2s_1}{v}+\frac{s_2}{v}"

"T=2\frac{20}{12.96}+\frac{100}{12.96}=10.8\ s"